LHC 2010 - latest news

Week 34: Machine coordinator: Ralph Assmann

Main aims of the week:

List of End of Fill studies (in work)

List of MD requests (in work)



Link to report from morning meeting

Monday 30th August

·        Stop of beam at 5h30 for technical stop.

Slides from 08h30 meeting

Sunday 29th August

·        Stable beams continuing, fill # 1308. ATLAS and CMS have now recorded more than 3,000 nb-1 at 3.5 TeV!

·        12h24: Programmed beam dump. Ramp down and pre-cycle.

·        13h45: Injection.

·        16h30: Start ramp.

·        18h17: Stable beams fill #1309.

Slides from 09h00 meeting

Saturday 28th August

·        02h23: Machine closed.

·        02h40: Switch A on RQTF.A56.B2 broken. Access for repair.

·        07h00: Machine closed and start of pre-cycle.

·        08h30: Injection of beam but staying at 450 GeV.

·        10h30: Decision to allow ramp to 3.5 TeV with modified UPS configuration. Regular physics operation only for the rest of the weekend. Broken UPS unit to be repaired on Monday during technical stop.

·        10h54: Start of ramp.

·        11h41: Flat top.

·        11h50: BLM triggered beam dump during chromaticity measurement. Losses with long integration time in warm quadrupole. Ramp down and pre-cycle.

·        13h19: Injection.

·        14h20: SPS scraper broken. Needs SPS access for repair. LHC beams partially filled, kept at 450 GeV.

·        19h25: SPS problem solved. Reinjecting LHC beams.

·        20h45: Start of ramp.

·        22h42: Stable beams fill #1308.

Slides from 9h00 meeting

Friday 27th August

·        01h38: Rampdown and precycle.

·        03h04: Injection for physics.

·        04h14: Start ramp.

·        06h11: Stable beams fill #1305. Delivered total luminosity reaches more than 3,000 nb-1.

·        09h41: Adjust mode: end of fill meas. of transverse tails (B2, H).

·        10h18: Ramp down and pre-cycle.

·        11h54: Bunch train commissioning (ADT, PLL, injection).

·        16h14: At time of injection loss of 4b train (4e11 p): lost QPS-OK on RB.A81, on magnet A8R8, close to transfer line collimator.

·        16h50: Alarm on UPS unit EBS21/28 in RE28. External company called in for diagnostics.

·        17h09: Lost QPS OK on RB.A78 :quench heaters discharging in A18R7 (HTS 4).

·        19h47: Beams lost. Sectors 12 and 23 lost with UPS problem. Access required.

·        20h00: Various accesses for UPS and QPS. UPS could not be repaired. Switching to other UPS (loss of redundancy).

·        23h24: Internal water fault on PC RB.A56.

Slides from 8h30 meeting

Thursday 26th August

·        00h00: Collimation piquet and CO controls expert working.

·        04h21: Stable beams fill #1303. Lumi > 1e31 in ATLAS and CMS.

·        17h35: Beams dumped with fast beam loss event.

·        17h40: Ramp down and precycle. Access in Pt 4 ( for QPS problem on worldFIP problem).

·        21h42: Injection of 1 nominal bunch per beam.

·        23h08: Ramp, squeeze, collide. Loss maps for collimation monitoring.

Slides from 8h30 meeting

Wednesday 25th August

·        06h30: BCT calibration.

·        08h00: Adjust mode for end-of-fill study. Scraping test for determining transverse beam profile (beam 2, vertical). As expected, beam dumped with BLM trigger when approaching beam core.

·        08h10: Ramp-down and pre-cycle.

·        09h02: Water leak error on cavity 7B1. Access. No real water leak, instrumentation problem.

·        09h51: Recovery. Cycle.

·        10h29: Lost QPS-OK on RB.A45: heater discharging. Access.

·        12h37: Recovery. Cycle.

·        14h14: Injection pilot. Starting transverse damper and PLL work.

·        14h39: PS booster problems. RF rephrasing problems. XPOC server update.

·        17h57: Bunch trains injected successfully. Transverse damper work.

·        19h30: SPS problems. Timing problem with sequencer: only pilot bunches available.

·        22h03: SPS timing problem fixed. Now collimator controls problem.

Slides from 8h30 meeting

Tuesday 24th August

·        00h11: Stable beams fill #1299.

·        03h27: Beams dumped due to RD1.R2 trip. Access by piquet.

·        06h23: Machine closed. Precycle started.

·        07h47: Problem with QPS OK on RB.A67 (worldFIP problem on the RB/RQD/RQF). Access for repair. Water problem with transverse damper. Access to P4. New transverse damper firmware for nominal bunch intensity (saturation problem).

·        11h05: Recovery from access. Trips of RSF2.A56B1 and RQTL8.L3B1. RF power: crowbar alarm on module 6B2 and 7B2.

·        13h02: Maintenance injetion lines (steering). Losses reduced by factor 2-3.

·        13h30: Injection into LHC. Injection phase +12 Hz on both beams. New injection scheme used: “1000ns_50b_35_14_35”.

·        15h39: Ramp started. Fill 1301.

·        17h35: Stable beams fill #1301. First time luminosity above 10e31!

Slides from 8h30 meeting

Monday 23rd  August

·        Continuing physics fill.

·        13h48: Beam dump due to fast beam losses in IR3 (MQ22.R3).

·        14h58: Accesses to point 4 (QPS server) and point 2 (MKI check).

·        17h43: Injection.

·        18h43: Trip of RD1.R2 (water) during injection process.  Reset by piquet.

·        19h38: Injection.

·        22h21: Start of ramp.

Slides from 8h30 meeting